OpenStudy (anonymous):

Can someone help me asap?! y = 6x-3 y = 8x-6

OpenStudy (anonymous):

What do you want the answer to?

OpenStudy (anonymous):

finding the solution, an ordered pair

OpenStudy (anonymous):

If I'm not completely wrong you substitute one of the letters for 0 at a time. So to solve for x you substitute y for 0 and vice versa...

OpenStudy (anonymous):

oh I think I get it.. like 6(0)-3 ?? im not sure..

OpenStudy (anonymous):

So, to solve for x: 0=6x-3 x=3/6 (1/3, 0) And to solve for y: y=6(0)-3 y=-3 (0, -3)

OpenStudy (anonymous):

The coordinates are the ordered pair. I'm not 100% sure, but I think that's how you do it.

OpenStudy (anonymous):

the answer says( 3/2 , 6 ) but idk how to do it >

OpenStudy (anonymous):

nevermind I got it you have to graph it (: x=1/2 and y = 6 Thanks ! :D

OpenStudy (anonymous):

Oh! That makes sense! Haha, I've just been sitting here staring at the page trying to figure it out, sorry! Good luck